\begin{thebibliography}{10}
\expandafter\ifx\csname url\endcsname\relax
  \def\url#1{\texttt{#1}}\fi
\expandafter\ifx\csname doi\endcsname\relax
  \def\doi#1{\burlalt{doi:#1}{http://dx.doi.org/#1}}\fi
\expandafter\ifx\csname urlprefix\endcsname\relax\def\urlprefix{URL }\fi
\expandafter\ifx\csname href\endcsname\relax
  \def\href#1#2{#2}\fi
\expandafter\ifx\csname burlalt\endcsname\relax
  \def\burlalt#1#2{\href{#2}{#1}}\fi

\bibitem{karlhede1980review}
Anders Karlhede.
\newblock A review of the geometrical equivalence of metrics in general
  relativity.
\newblock {\em General Relativity and Gravitation}, 12:693--707, 1980.

\bibitem{kramer}
H.~Stephani, D~Kramer, M.~MacCallum, C.~Hoenselaers, and E.~Herlt.
\newblock {\em Exact Solutions of Einstein's Field Equations}.
\newblock Cambridge University Press, 2009.

\bibitem{mcnutt2017cartan}
D.~D. McNutt, A.~A. Coley, and A.~Forget.
\newblock The cartan algorithm in five dimensions.
\newblock {\em Journal of Mathematical Physics}, 58(3), 2017.

\bibitem{eichhorn2021locality}
A.~Eichhorn and A.~Held.
\newblock From a locality-principle for new physics to image features of
  regular spinning black holes with disks.
\newblock {\em Journal of Cosmology and Astroparticle Physics}, 2021(05):073,
  2021.

\bibitem{kruglikov2022killing}
B.~Kruglikov and W.~Steneker.
\newblock Killing tensors in koutras--mcintosh spacetimes.
\newblock {\em Classical and Quantum Gravity}, 39(22):225005, 2022.

\bibitem{Kitamura1978}
{Kitamura, S.}
\newblock {A C.S. method for stationary axially symmetric space-times and an
  application to generalizations of G{\"o}del's universe}.
\newblock {\em Tensor N.S.}, 32:156, 1978.

\bibitem{marvan2008local}
M.~Marvan and O.~Stolin.
\newblock On local equivalence problem of space-times with two orthogonally
  transitive commuting killing fields.
\newblock {\em Journal of mathematical physics}, 49(2):022503, 2008.

\bibitem{ferraioli2020equivalence}
D.~C. Ferraioli and M.~Marvan.
\newblock The equivalence problem for generic four-dimensional metrics with two
  commuting killing vectors.
\newblock {\em Annali di Matematica Pura ed Applicata (1923-)},
  199(4):1343--1380, 2020.

\bibitem{robinson1975uniqueness}
D.~C. Robinson.
\newblock Uniqueness of the kerr black hole.
\newblock {\em Physical Review Letters}, 34(14):905, 1975.

\bibitem{mars2000uniqueness}
M.~Mars.
\newblock Uniqueness properties of the kerr metric.
\newblock {\em Classical and Quantum Gravity}, 17(16):3353, 2000.

\bibitem{ferrando2009intrinsic}
J.~J. Ferrando and J.~A. S{\'a}ez.
\newblock An intrinsic characterization of the kerr metric.
\newblock {\em Classical and Quantum Gravity}, 26(7):075013, 2009.

\bibitem{Owen:2021eez}
C.~B. Owen, N.~Yunes, and H.~Witek.
\newblock {Petrov type, principal null directions, and Killing tensors of
  slowly rotating black holes in quadratic gravity}.
\newblock {\em Physical Review D}, 103(12):124057, 2021.
\newblock \href{http://arxiv.org/abs/gr-qc/2103.15891}{{arXiv:2103.15891
  [gr-qc]}}.

\bibitem{schwarzschild1916gravitationsfeld}
K.~Schwarzschild.
\newblock {\"U}ber das gravitationsfeld eines massenpunktes nach der
  einsteinschen theorie.
\newblock {\em Sitzungsberichte der k{\"o}niglich preussischen Akademie der
  Wissenschaften}, pages 189--196, 1916.

\bibitem{cahen1968lorentzian}
M.~Cahen and L.~Defrise.
\newblock Lorentzian 4 dimensional manifolds with “local isotropy”.
\newblock {\em Communications in Mathematical Physics}, 11:56--76, 1968.

\bibitem{Wald:1984rg}
R.~M. Wald.
\newblock {\em {General Relativity}}.
\newblock Chicago University Press, 1984.

\bibitem{greenberg1970propagation}
P.~J. Greenberg.
\newblock Propagation equations for the set of natural geometrical invariants
  in relativistic hydrodynamics.
\newblock {\em Journal of Mathematical Analysis and Applications},
  29(3):647--667, 1970.

\bibitem{chinea1992differential}
F.~J. Chinea and L.~M. Gonz{\'a}lez-Romero.
\newblock A differential form approach for rotating perfect fluids in general
  relativity.
\newblock {\em Classical and Quantum Gravity}, 9(5):1271, 1992.

\bibitem{fernandez1999exterior}
L.~Fern{\'a}ndez-Jambrina and L.~M. Gonz{\'a}lez-Romero.
\newblock Exterior differential system for cosmological \protect{G}2 perfect
  fluids and geodesic completeness.
\newblock {\em Classical and Quantum Gravity}, 16(3):953, 1999.

\bibitem{fayos1999papapetrou}
F.~Fayos and C.~F. Sopuerta.
\newblock On the papapetrou field in vacuum.
\newblock {\em Classical and Quantum Gravity}, 16(9):2965, 1999.

\bibitem{fayos2001general}
F.~Fayos and C.~F. Sopuerta.
\newblock General approach to the study of vacuum spacetimes with an isometry.
\newblock {\em Classical and Quantum Gravity}, 18(3):353, 2001.

\bibitem{fayos2002consequences}
F.~Fayos and C.~F. Sopuerta.
\newblock Consequences of a killing symmetry in spacetime's local structure.
\newblock {\em Classical and Quantum Gravity}, 19(21):5489, 2002.

\bibitem{katona2023classification}
D.~Katona.
\newblock A classification of supersymmetric kaluza-klein black holes with a
  single axial symmetry.
\newblock {\em arXiv preprint arXiv:2306.09933}, 2023.

\bibitem{katona2023supersymmetric}
D.~Katona and J.~Lucietti.
\newblock Supersymmetric black holes with a single axial symmetry in five
  dimensions.
\newblock {\em Communications in Mathematical Physics}, 399(2):1151--1201,
  2023.

\bibitem{mcnutt2023frame}
D.~D. McNutt, A.~A. Coley, and R.~J. van~den Hoogen.
\newblock A frame based approach to computing symmetries with non-trivial
  isotropy groups.
\newblock {\em Journal of Mathematical Physics}, 64(3), 2023.

\bibitem{olver1995equivalence}
P.~J. Olver.
\newblock {\em Equivalence, invariants and symmetry}.
\newblock Cambridge University Press, 1995.

\bibitem{chinea1988symmetries}
F.~J. Chinea.
\newblock Symmetries in tetrad theories.
\newblock {\em Classical and Quantum Gravity}, 5(1):135, 1988.

\bibitem{estabrook1996moving}
F.~B. Estabrook and H.~D. Wahlquist.
\newblock Moving frame formulations of 4-geometries having isometries.
\newblock {\em Classical and Quantum Gravity}, 13(6):1333, 1996.

\bibitem{papadopoulos2012locally}
G.~O. Papadopoulos and Th. Grammenos.
\newblock Locally homogeneous spaces, induced killing vector fields and
  applications to bianchi prototypes.
\newblock {\em Journal of mathematical physics}, 53(7), 2012.

\bibitem{garcia2009spinor}
A.~Garc{\'\i}a-Parrado G{\'o}mez-Lobo and J.~M. Mart{\'\i}n-Garc{\'\i}a.
\newblock Spinor calculus on five-dimensional spacetimes.
\newblock {\em Journal of mathematical physics}, 50(12), 2009.

\bibitem{pugliese2022wormholes}
D.~Pugliese and H.~Quevedo.
\newblock Wormholes, killing horizons and naked singularities: light surfaces
  in axially symmetric spacetimes.
\newblock {\em The European Physical Journal C}, 82(12):1090, 2022.

\end{thebibliography}

\documentclass[12pt,letterpaper]{article}
\usepackage[utf8]{inputenc}
\usepackage[english]{babel}
\usepackage[vmargin=0.6in,hmargin={1in,1in},headsep=0.3in,headheight=1in]{geometry}
\usepackage{subfiles}
\usepackage{tikz}
\usepackage{fix-cm}
\usepackage{enumitem}
\usepackage{multicol}
\usepackage{amsmath, amsthm, amssymb, amsfonts,mathrsfs}
\usepackage{paracol}
\usepackage{graphicx}

\allowdisplaybreaks
\newcounter{taggedEquations}
\let\OldTag\tag
\renewcommand*{\tag}[1]{\stepcounter{taggedEquations}\OldTag{#1}}
\usepackage{array}
\usepackage{tabularx}

\usepackage{relsize}
\usepackage{xcolor}
\usepackage{lastpage}
\usepackage{blindtext}
\usepackage[numbers,sort&compress]{natbib}
\usepackage[linktocpage=true]{hyperref}
\hypersetup{
	colorlinks,
	citecolor=blue,
	filecolor=black,
	linkcolor=blue,
	urlcolor=blue
}
\usepackage{url}

\usepackage{graphicx}
\usepackage{caption}
\usepackage{subcaption}
\usepackage{wrapfig}
\usepackage{indentfirst}
\usepackage{hyperref}
\usepackage{soul}
\usepackage{bm}

\usepackage{empheq}
\usepackage{authblk}
\setlist{parsep=0pt,listparindent=\parindent}
\setlist[itemize]{noitemsep, topsep=0pt}
\setlist[enumerate]{noitemsep, topsep=0pt}
\setlist{parsep=0pt,listparindent=\parindent}

\stepcounter{footnote}
\graphicspath{{Images/}{../Images/}}
\usepackage{graphicx}
\graphicspath{ {./images/} }
\usepackage{pgfplots}
\usepackage{extarrows}
\usepackage{mathtools}
\usepackage{tikz}
\usetikzlibrary{decorations.pathmorphing,patterns}
\usetikzlibrary{decorations.pathreplacing,angles,quotes}
\usetikzlibrary{patterns}
\usetikzlibrary{shapes.arrows, fadings}
\usepackage{xfrac} 
\usepackage{relsize}
\usepackage[most]{tcolorbox}
\tikzfading[name=fade down,
  top color=transparent!0, bottom color=transparent!100]
\tikzfading[name=fade up,
  bottom color=transparent!0, top color=transparent!100]

\usepackage{pgfplots}
\pgfplotsset{compat=1.18}
\usetikzlibrary{external}

\usepackage{physics}

\usepackage{enumitem}

\def\beq{\begin{eqnarray}}  
\def\eeq{\end{eqnarray}}

\begin{document}

\title{\Large\textbf{Killing Invariants: An approach to the sub-classification of geometries with symmetry}}
\author[1,2]{\large{Brown, C. }}
\author[1,2]{\large{Gorban, M.}}
\author[1,2]{\large{Julius, W.}}
\author[1,2]{\large{ Radhakrishnan, R.}}
\author[1,2]{\large{Cleaver, G. }}
\author[3]{\large{McNutt, D. }}

\affil[1]{\emph{Early Universe, Cosmology and Strings (EUCOS) Group, Center for Astrophysics, Space Physics and Engineering Research (CASPER)}}
\vspace{1 cm}
\affil[2]{\emph{Department of Physics, Baylor University,  Waco, TX 76798, USA} }
\vspace{1 cm}
\affil[3]{\emph{UiT: The Arctic University of Norway, Department of Mathematics and Statistics, 9019, Tromsø, Norway, }}
\vspace{0.25cm}

\date{\today}  
\maketitle  
\begin{abstract}   

In principle, the local classification of spacetimes is always possible using the Cartan-Karlhede algorithm. However, in practice, the process of determining equivalence of two spacetimes is potentially computationally difficult or not at all possible. This difficulty will arise whenever the classifying invariants are either high-degree rational functions or depend on transcendental functions without standard inverses. In the case that spacetimes admit Killing vectors with non-trivial orbits, we propose a new set of invariant quantities, called Killing invariants. These invariants will allow for the sub-classification of spacetimes admitting the same group of symmetries and will, in principle, be substantially less complicated than any other known set. We apply this approach to the class of static spherically symmetric geometries as an illustrative example. 

\end{abstract} 

\par

\pagebreak

\section{Introduction}\label{section:intro}

The Cartan-Karlhede (CK) algorithm can always be applied to locally classify Lorentzian geometries  \cite{karlhede1980review}. It has been used to distinguish solutions within General Relativity (GR) \cite{kramer} and alternative metric-based gravity theories in four dimensions (4D) and higher dimensions \cite{mcnutt2017cartan}. In principle, the algorithm provides a method to determine when two line-elements are equivalent; that is, when a coordinate transformation can be used to transform one line-element into the other. However, in practice, the algorithm is primarily used to prove when spacetimes are inequivalent, and hence distinct. This is due to the final step used to compare spacetimes, where one must equate the corresponding sets of Cartan invariants of each line-element and show that these resulting equations have a solution. This step is not necessarily algorithmic and may result in an unsolvable system of equations if high-order rational functions or transcendental functions appear in these expressions. 

An excellent example of this difficulty appears within the class of stationary axisymmetric spacetimes with two commuting Killing vector fields \cite{kramer}. If a given spacetime admits an irreducible rank 2 Killing tensor field, then the line-elements within this class of spacetimes  can be expressed in Boyer-Lindquist form  \cite{eichhorn2021locality}. However, determining the existence of this tensor can be difficult \cite{kruglikov2022killing}. Given a stationary axisymmetric metric in an arbitrary coordinate system and one in Boyer-Lindquist coordinates, it is then possible to determine if they are equivalent. The inequivalence can be determined by comparing the discrete invariants generated in the algorithm, namely the number of functionally independent invariants and the dimension of a subgroup of the Lorentz frame transformations determined at each iteration of the algorithm. If two line-elements share the same discrete invariants, then the final step must be initiated to show equivalence. For these spacetimes, the coordinate expressions of the Cartan invariants are quite complicated and the resulting equations are typically difficult to solve. 

To address this, other approaches have been used, either by studying the kinematic properties of a chosen frame basis for the spacetimes 
\cite{Kitamura1978} or by working with differential invariants associated with the jet bundle \cite{marvan2008local, ferraioli2020equivalence}. In the latter two papers, the authors'  claim that the coordinate expressions for the differential invariants are less complicated than those of the Cartan invariants. While this is true, the resulting invariants are still very large and the problem of determining equivalence is not substaintially easier than before. 

This problem is of particular importance when exploring the class of axisymmetric and stationary vacuum solutions in GR beyond the Kerr solution. While there are uniqueness  theorems and classification results for the Kerr black hole \cite{robinson1975uniqueness, mars2000uniqueness, ferrando2009intrinsic}, the investigation of other stationary axisymmetric solutions within GR or alternative gravity theories \cite{Owen:2021eez, eichhorn2021locality} will require practical tools to classify solutions. 

The issue of complicated coordinate expressions for curvature invariants can be alleviated by introducing a new set of invariants for a spacetime, whose coordinate expressions are expected to be smaller than any set of invariants used previously. To construct these invariants, we will consider quantities generated from Killing vector fields in the spacetime. If a spacetime admits Killing vector fields, these are certainly invariantly defined up to scaling by constant values and linear combinations. A natural question to ask is whether or not scalar invariants can be constructed using them. 

Geometrically, the use of Killing vector fields as a tool to generate solutions is well-known \cite{schwarzschild1916gravitationsfeld,cahen1968lorentzian,kramer}. However, the perspective that Killing vector fields have some relevance to physics is less obvious. The existence of symmetries in a spacetime implies the existence of conserved quantities when studying geodesics \cite{Wald:1984rg}. More generally, it has been argued that the kinematics of spacelike or timelike geodesics that lie exclusively in the orbits of the symmetry group give rise to physically meaningful invariants \cite{ greenberg1970propagation, chinea1992differential, fernandez1999exterior}.

Fayos and Sopuerta \cite{fayos1999papapetrou, fayos2001general} detail a method of studying the local structure of spacetime by using the Papapetrou field, the exterior derivative of the Killing vector field in their definition, as the main algebraic object. They emphasize the use of the Killing vector field as a vector potential solution of Maxwell's equations and the Papapetrou field therefore being seen as an electromagnetic field tensor. Through the integrability conditions of the Killing vectors, written in terms of the Papapetrou fields, they show that the Weyl tensor in 4D can be completely expressed in terms of the Killing vector field, the Papapetrou field, and the energy-momentum tensor. In addition, by algebraically classifying Papapetrou forms, they determine conditions for when the principal directions of a Papapetrou field aligned with the principal directions of the Weyl tensor and hence determine the Petrov type of the spacetime \cite{fayos1999papapetrou,fayos2001general,fayos2002consequences}. It is also shown that the Newman-Penrose formalism can be adapted into a form that places the Papapetrou field into a canonical form. These results on Papapetrou fields can also be extended by adding new variables derived from the Killing vectors. Through this, other values such as the Ricci tensor components can be algebraically expressed in terms of the matter fields \cite{fayos2002consequences}. These results indicate that the study of Killing vector fields in general relativity can give meaningful information about the physical interpretation of solutions.

The use of invariant quantities derived from Killing invariants is not restricted to 4D GR or its modifications. For example, the classification of five-dimensional supersymmetric black holes with a single axial symmetry was accomplished using the norm of a Killing vector field as a global invariant for the solutions \cite{katona2023classification, katona2023supersymmetric}. In gravitational theories involving torsion, such as Einstein-Cartan or teleparallel gravity, a frame formalism can be adapted to the group of symmetries in order to derive the most general family of solutions that admit that symmetry group. These invariants also provide a direct way to sub-classify these families of solutions \cite{mcnutt2023frame}.

We will adopt a frame approach similar to the CK algorithm, although the algorithm will also share features of Cartan's method of moving frames \cite{olver1995equivalence}. We will use the fact that the Killing vector fields of a spacetime are invariantly defined and employ them to fix the frame in a coordinate independent manner by treating the Killing one-forms relative to a chosen coframe as an exterior differential system \cite{chinea1988symmetries, estabrook1996moving, papadopoulos2012locally}. This approach will take a given Lie algebra of Killing vector fields and generate a representative frame basis for the manifold. While the stationary axisymmetric spacetimes have been discussed in order to motivate this algorithm, we will instead focus on the class of spherically symmetric spacetimes in order to introduce our formalism, and in a future paper we will apply this approach to stationary axisymmetric geometries.

\subsection{Generating Killing vector fields}

As a partial answer to the question of scaling by a constant value, we can employ the Cartan-Karlhede algorithm to determine an invariantly defined frame, up to freedom in the linear isotropy group. We will call this frame a {\it CK frame} and write this as $\{{\bf e}_a\}$ and the dual coframe as $\{ \theta^a\}$. In the case of non-trivial isotropy, the pullback of a diffeomorphism, $\Phi: M \to M$, gives 
\beq
    \Phi_* \theta^a = \Lambda^a_{~b} \theta^b, 
\eeq
\noindent where $\Lambda^a_{~b}$ lies in the linear isotropy group $H_p(M)$.  If $\Phi_\tau$ is the flow of a Killing vector field, ${\bf X}$, at a point $p \in  M$, then taking the limit of $\tau\to 0$ we find 
\beq 
    \mathcal{L}_{{\bf X}} \theta^a = \lambda^a_{~b} \theta^b, 
\eeq

\noindent where $\lambda^a_{~b}$ lies within the Lie algebra of the linear isotropy group. For a given CK frame, we can solve the resulting frame based Killing equations to determine the form of the Killing vector fields \cite{mcnutt2023frame}. In doing so, we can write any Killing vector field, ${\bf X}$, of a given spacetime in terms a CK frame where the coefficients, $X^a$, are invariant quantities, 
\beq 
    {\bf X} = X^a {\bf e}_a.
\eeq
\noindent The Killing vector-fields are solutions of a linear first order partial differential equation, the Killing equations, and so it is useful to determine a basis of solutions. Relative to the CK frame, we can determine the set of Killing vector fields in a coordinate independent manner by identifying a linear independent basis of solutions to the Killing equations. We can effectively ignore the problem of scaling the Killing vector fields so long as an invariantly defined frame has been constructed and we are only comparing functional relationships between the resulting invariants. 

The CK frame does not resolve the ambiguity involved in constant linear combinations of Killing vector fields. For example, in the class of stationary axisymmetric spacetimes, we can apply a $GL(2, \mathbb{R})$ transformation to produce new timelike and spacelike Killing vectors which cannot be distinguished from the original pair. This is not unexpected nor unwelcome, it is a natural byproduct of the symmetry of the spacetime. We note that the existence of a non-trivial Lie algebra can restrict the freedom of linear combinations by imposing a canonical form for the structure coefficients. 

We will consider the case of an $N$-dimensional manifold. The procedure begins with the assumption that the set of all Killing vector fields, $\{ {\bf X}_A \}$, $A,B,E = 1,\ldots, m$ with $m \leq \frac{N(N+1)}{2}$  is known in the coordinate basis. We will denote the Killing one-forms as ${\bf F}^A = g_{ab} X^a_{~A} \theta^b$. The structure coefficients $C^A_{~BE}$ and inner-products of the Killing vector fields, 
\beq  N_{AB} = g_{ab} X^{\,a}_{~A} X^{\,b}_{~B},  \eeq
are then computed to determine where the vector fields are timelike, spacelike, or null. We will call the inner-product of two copies of a vector-field the norm.  We then determine if there are points in the manifold where the norm of the Killing vector-field vanishes, $N_{AA} = 0 $, and $ X^a_{~A} = 0$. If this occurs for a given Killing vector it lies in the isotropy group. If a given Killing vector is not a member of the isotropy group it is instead a translational Killing vector, which we  collect and form the set  $\{ {\bf {\bf X}_I}\}$, $I, J, K = 1, \ldots m'$ where $m' \leq r$ and $r \leq N$ is the dimension of the orbits of the symmetry group. For example, in the case of spherical symmetry in N dimensions, the orbits of the symmetry group is $m' = N-1$ whereas the translational Killing vector fields is at most $2$ in a coordinate system where one spatial rotation has been rectified. 


Taking the perspective that the $N_{IJ}$ have geometrical and physical significance, the vanishing of the norm, $N_{II}$, indicates where the translational Killing vectors become null. For example, in the exterior of the ergosphere of the Kerr black hole, the timelike Killing vector field, $\partial_t$, becomes null at any point in the ergosphere. On surfaces where $N_{II}$ vanish, it is not possible to construct an orthonormal frame and, thus, we must restrict the neighbourhood appropriately to work with an orthonormal frame. In the context of classification, this is not an obstacle and potentially provides a helpful diagnostic to check the inequivalence of two solutions.

In an open region where all $N_{II}$ are non-zero, the Gram-Schmidt procedure constructs an orthonormal frame adapted to the translational Killing vector fields. As an initial coframe, one could use the simplest coframe arising from the metric or the CK coframe. The result of this process is a coframe, $\{ \theta^a \}$, and a restricted subgroup of Lorentz frame transformations $H_{-1}$. We will denote the dual frame basis as $\{ \mathbf{e}_a \}$. Relative to this frame, the components of the translational Killing vector fields will be written as rational expressions of the $N_{IJ}$.  

The algorithm to construct invariant quantities from a set of Killing vector fields is then:
\begin{enumerate}
\item Denote the set $\mathcal{K}_0 =  \{ N_{IJ} \} $ for $I, J = 1, \ldots, m'$ as the zeroth-order Killing invariants.
\item Set the order of differentiation q to one. 
\item Calculate the exterior derivatives of the Killing one-forms, $d{\bf F}^I = F^I_{~ab} \theta^a \wedge \theta^b$ where $F^I_{~ab} = e^{}_{[a}(F^I_{~b]})$ where $e_{a_i}$ acts as a frame derivative operator on a scalar.
\item Find the canonical form for each of the $d{\bf F}^I$.
\item Fix the frame using this canonical form and note the residual frame freedom, which we will call the linear isotropy group, $\bar{H}_0$. 
\vspace{2 mm}

\noindent The components of $d{\bf F}^I$ relative to this frame are called first-order Killing invariants, denoted as an ordered list \beq 
    \mathcal{K}_1 = \mathcal{K}_0 \cup \{ F^I_{~ab} \}, \quad\quad 0 \leq a<b \leq 4. 
\eeq 
Similarly, we will introduce the ordered list $\mathcal{K}_q$ as the collection of all Killing invariants up to $q^{th}$-order, 

\beq 
    \mathcal{K}_q =  \mathcal{K}_{q-1} \cup \{ F^I_{~b a_1 a_2 \ldots a_q} \},
\eeq

\noindent where $ F^I_{~b a_1 a_2 \ldots a_q} = - e_{a_q} \ldots e_{a_2} e^{}_{[a_1}(F^I_{~b]})$.
\vspace{2 mm}

\item For $q\geq 1$ calculate the exterior derivatives of the elements in $\mathcal{K}_q$. 
\item Fix the frame as far as possible using the one-form, $d k$ with $k \in \mathcal{K}_q$, and note the residual frame freedom, which we will call the linear isotropy group, $\bar{H}_q$. 

\item If the linear isotropy group is the same as the previous iteration, let $p = q$ and stop; if they differ, or if $q=0$, increment $q$ by 1 and go to step 5.  
\end{enumerate}

\noindent Upon completing the algorithm we will denote $\mathcal{K} = \mathcal{K}_{p}$ as the set of all Killing invariants. Any spacetime admitting a symmetry group can then be characterized locally by the canonical form used and the successive isotropy groups. The Killing invariants in $\mathcal{K}$ can then be used to characterize the geometry.

To compare two spacetimes, $(M,{\bf g})$ and $(M', {\bf g'})$, we can use the descending chain of linear isotropy groups at each order to determine inequivalence. We note that comparing their respective sequences is not sufficient to determine equivalence. If the two sequences agree so that the spacetimes share similar canonical forms, it is necessary to compare the set of equations obtained by equating each entry in the ordered list $\mathcal{K}_{p}$ with the corresponding entry in $\mathcal{K}'_{p}$. If this set of equations has a solution, then the two spacetimes are equivalent. We note that this final step is not algorithmic and is potentially unsolvable. However, in practice, this is often achievable.


\section{Static spherically symmetric metrics} \label{section:SSSmetrics}

To illustrate how the algorithm can be implemented when non-trivial isotropy is present, we consider the class of N-dimensional spherically symmetric metrics,
\beq ds^2 = -e^{2\alpha(r)} dt^2 + e^{2\beta(r)}dr^2 + r^2 d\Omega_N, \label{eq:StandardSSM_HD} \eeq
\noindent where 
\beq d \Omega_N = d\phi_1 d \phi_1 + \sum_{\mu = 2}^{N} S^2_{\mu} d \phi_\mu d \phi_\mu, \label{eq:Nsphere} \eeq
\noindent and 
\beq S_{\mu}(\phi_1, \ldots \phi_{\mu-1}) = \left( \prod_{\nu = 1}^{\mu-1}  \sin (\phi_\nu) \right). \eeq

The Lie algebra of Killing vector fields is isomorphic to $\mathbb{R} \times \mathfrak{so}(N+1)$. In Cartesian coordinates, the vector fields lying in $\mathfrak{so}(N+1)$ vanish at some point and hence lie in the isotropy group. However, in these coordinates, we have two translational Killing vector fields,
\beq {\bf X}_1 = \partial_t, \quad\quad {\bf X}_2 = \partial_{\theta_N}, \eeq
\noindent or as one-forms: 
\beq {\bf F}^1 = e^{2\alpha} dt, \quad\quad{\bf F}^{2} = r^2S_N^2 d \phi_N. \eeq

\noindent The second one-form is well-defined except when $\theta^N =0$. With these vector fields, we may choose them as frame elements almost everywhere in the geometry.

In standard coordinates, the Gram-Schmidt procedure yields the following basis written as a null coframe, $\{\theta^a\} = \{ n,\ell, \bar{m}, m, m_\mu\}$, \cite{kramer,garcia2009spinor}
\beq \begin{aligned}
n & = \frac{1}{\sqrt{2}}( e^\alpha dt - r S_N d\phi_N), \\[10pt]
\ell & = \frac{1}{\sqrt{2}}( e^\alpha dt + r S_N d\phi_N), \\[10pt]
\bar{m} & = \frac{1}{\sqrt{2}}( e^\beta dr + i r d \phi_1), \\[10pt]
m &= \frac{1}{\sqrt{2}}( e^\beta dr - i r d\phi_1 ), \\[10pt]
m_\mu &= r S_\mu d \phi_\mu, \quad\quad \mu = 2,\ldots, N-1 .
\end{aligned} \eeq

At zeroth-order, we find two non-trivial Killing invariants, 
\beq N_{11} = e^{2\alpha}, \quad\quad N_{22} = r S_N .\eeq

Taking the exterior derivatives of ${\bf F}^1$ and ${\bf F}^2$, we find
\beq \begin{aligned} d {\bf F}^1 &= -4 \alpha_{,r} e^{\alpha-\beta} (m + \bar{m}) \wedge (n+\ell), \\
d {\bf F}^2 &= -e^{-\beta} S_N (m+\bar{m}) \wedge (n-\ell ) - i S_N \frac{\cos(\phi_1)}{\sin(\phi_1)}  (m - \bar{m}) \wedge (n - \ell) \\
&~~~- \sum_{\mu = 2}^{N-1} \frac{\sqrt{2} S_N }{S_{\mu}} \frac{\cos(\phi_\mu)}{\sin(\phi_\mu)} [m^{\mu} \wedge (\ell-n)].  \end{aligned} \eeq

\noindent The remaining parameters for the boost and spatial rotations can be set to identity, as the components of the two-forms are described entirely by their squared-norms. To conclude the algorithm, we would compute the exterior derivatives of the components $\{ F^I_{~ab}\}$, as the linear isotropy group would remain trivial and indicate the algorithm stops at order $2$. Thus $\mathcal{K}_2$ contains all relevant information and spherically symmetric metrics can be classified using the classifying functions arising from this set. 

\subsection{Static spherically symmetric metrics with a Killing horizon}

A Killing horizon is defined as a null hypersurface where a Killing vector becomes null and acts a generator for the null hypersurface \cite{Wald:1984rg}. On such a surface, it is no longer possible to build an orthonormal frame using the translational Killing vector fields. However, Killing invariants can be applied to any region outside of this surface. 

In any dimension, such a metric takes the following form in advanced coordinates
\beq ds^2 = - e^{2 \gamma(r)} \left( 1 - \frac{2m(r)}{r}\right) dv^2 +2 dvdr + r^2 d \Omega_N, \eeq

\noindent where $d\Omega_N$ is defined in Equation \eqref{eq:Nsphere}. In this coordinate system the timelike Killing vector is given as $ {\bf X}_1 = \partial_v$. Using the transformation $t = v + f(r)$, where the function $f$ satisfies $f'(r) = \frac12 e^{-\gamma(r)} \left(1-\frac{2m(r)}{r}\right)^{-1}$, the metric becomes 
\beq ds^2 = - e^{2\gamma(r)} \left( 1 - \frac{2 m(r)}{r}\right) dt^2 + \left( 1 - \frac{2 m(r) }{r}\right)^{-1} dr^2 + r^2 d\Omega_N. \eeq

\noindent Comparing with the general metric in Equation \eqref{eq:StandardSSM_HD}, it follows that 
\beq \alpha(r) = \gamma(r) + \frac12 \ln \left( 1 - \frac{2m(r)}{r}\right), \quad\quad \beta(r) = - \frac12 \ln \left( 1 - \frac{2m(r)}{r}\right), \eeq
\noindent while two relevant Killing invariants take the form
\beq \begin{aligned} N_{11} &= e^{2 \gamma(r)} \left( 1 - \frac{2m(r)}{r}\right), \\
F^1_{~13} &= \frac{e^{\gamma(r)}}{r^2} \qty[ r \gamma(r)_{,r}  (2 m(r) - r) + rm(r)_{,r} - m(r)]. \end{aligned} \eeq

Given a scalar invariant, its zero-set specifies an invariantly-defined hypersurface, which can have physical significance for the spacetime. In the current expressions we may determine coordinate values for the Killing horizon as their zero-set. In general, the second invariant, $F^1_{~13}$, will not vanish on the horizon, $r=2m$, except for special choices of $m(r)$; thus, we can choose $F^1_{~13}$ as the functionally independent invariant. Using the inverse function theorem, we may express the first invariant in terms of the other, which we will call a classifying curve for the space as this expression will be independent of coordinates:
\beq I_{r} \equiv F^1_{~13}; \quad\quad N_{11} = e^{2\gamma'(I_r)} \left( 1 - \frac{2m'(I_r)}{r(I_r)} \right). \eeq 
\noindent If both of the invariants vanish on the Killing horizon, due to the choice of $m(r)$, the ratio of these two Killing invariants is again a Killing invariant, which is non-zero at all points in the manifold. We may then write down a different classifying curve for the space:
\beq I_r \equiv \frac{N_{11}}{F^1_{~13}};\quad\quad  N_{11} = \sqrt{I_r}\left( 1 - \frac{2m'(I_r)}{r(I_r)} \right), \quad\quad F^1_{~13} = \frac{1}{\sqrt{I_r}} \left( 1 - \frac{2m'(I_r)}{r(I_r)} \right). \eeq

\noindent Thus the Killing horizon can be expressed in an invariant manner, independent of any coordinate system. Furthermore, this can be generalized to solutions that may contain multiple Killing horizons, such as Schwarzschild(-anti)-de Sitter solutions or Reissner-Nordstr{\"o}m solutions.  

\section{Discussion}

In this paper, we have introduced a new formalism and set of invariants for classifying spacetimes admitting isometries. Any isometry of a space is generated by a Killing vector field, which is necessarily an invariant vector field of the spacetime and can be found by solving the Killing equations. Furthermore, the norm of any vector-field is an invariant quantity which is independent of the coordinate system chosen. By adapting a frame to the set of translational Killing vector fields, it is possible to work with their norms and exterior derivatives as an exterior differential system and generate a finite list of invariants. 

While determination of an invariant frame is a common tool in both the Cartan-Karlhede (CK) algorithm \cite{kramer} and Cartan's method of moving frames \cite{olver1995equivalence}, in general this approach is distinct from them. Only in the case of locally homogeneous or maximally symmetric spaces does this approach reduce to the Cartan's method of moving frames. However, our algorithm will never specialize to the CK algorithm since it does not use the irreducible parts of the curvature tensor or their covariant derivatives. Furthermore, the linear isotropy group of the CK algorithm will generically be larger than the linear isotropy group arising in our algorithm. 

In future work, we will employ the algorithm to characterize axisymmetric and stationary solutions to Einstein's field equations and other modified gravity theories in order to characterize these solutions and identify the subset of spacetimes which admit Boyer-Lindquist coordinates. Of particular interest are deformations of Kerr-like solutions \cite{pugliese2022wormholes} and whether such deformations preserve the necessary Killing tensor \cite{eichhorn2021locality}. 

\section*{Acknowledgments}
\noindent The authors would like to thank Eivind Schneider for their discussion and helpful comments. DM is supported by the Norwegian Financial Mechanism 2014-2021 (project registration number 2019/34/H/ST1/00636).

\bibliographystyle{unsrt-phys-eucos}
\bibliography{ref}

\end{document}